# The Bayesian Origin of the Probability Weighting Function in Human Representation of Probabilities

### v1

@@ 4 Empirical Validation of the Bayesian Framework
We empirically evaluate our theory within a unified Bayesian framework that allows direct head-to-head comparison between different model components (priors and encodings), as well as with the BLO model (Zhang et al., 2020). First, we validate Predictions 1, 3, 4 using a judgment of relative frequency (JRF) task, where subjects estimate the proportion of dots. Second, we assess the generality of our framework by applying it to a different domain, decision-making under risk. Third, we test our model’s ability to adapt to different stimulus statistics to test Prediction 2.
@@ 4.1 Validating Predictions on Judgment of Relative Frequency (JRF) data
We analyze data from Zhang et al. (2020) and Zhang & Maloney (2012), where subjects ($N=86$) judge the percentage of dots of a target color in arrays of black and white dots on a gray background. [^1] A complete description of the tasks and datasets used in this paper is available in Appendix C.1.
For the Bayesian model [^2] Additional model variants and detailed experimental results are available in Appendix C.6; detailed results for the other tasks discussed in Sections 4.2 and 4.3 also be found in the appendix., we evaluated two priors: (i) Gaussian prior with fitted mean and variance (GaussianP), and (ii) nonparametrically fitted prior (FreeP). For the encoding $F(p)$, we evaluated: (i) uniform encoding, $F(p)=p$ (UniformE), (ii) bounded-log-odds encoding $F(p)=\log(p+\beta_{1})-\log(1-p+\beta_{2})$ [^3] As in (11) but with separate upper and lower bounds, in line with BLO’s $\Delta_{-},\Delta_{+}$ bounds. (BoundedLOE), and (iii) freely fitted $F$ (FreeE). We also included a matched prior–encoding model to test the efficient coding account. In both BLO and the Bayesian model, parameters are optimized for each subject individually by maximizing trial-by-trial data likelihood. We fitted the Bayesian model using the method of Hahn & Wei (2024), which permits identifying the encoding nonparametrically from behavioral responses (Hahn et al., 2025). [^4] Across all datasets, we ran all models using the publicly available implementation of Hahn & Wei (2024), using a grid of 200 points and regularization strength 1.0.
@@
Figure 4: Resources of Bayesian FreeE FreeP model. IQR denotes interquartile range across subjects.
Supporting Prediction 1, we find two converging pieces of evidence. First, models with nonuniform encoding capture the characteristic S-shaped bias in behavior, whereas the uniform encoding fails to do so (Figure 3A, top). Second, the corresponding resources are U-shaped with peaks near the boundaries: the nonparametrically fitted resources show this pattern(Figure 4), qualitatively resembling the inherently U-shaped bounded log-odds encoding. The BLO model also shares the same overall U-shaped resources, but with two discontinuities—a property proved in Appendix Theorem 8 and illustrated in Appendix Figure 11. Moreover, recovered resources and prior show distinct shapes (Appendix, Figure 33); unlike predicted by efficient coding (Frydman & Jin, 2023), matching prior and encoding achieved substantially inferior fit compared to decoupled fitting of both components (Appendix, Figure 8).
In line with Prediction 3, the bimodal pattern of response variability with a dip near 0.5 is only explained by models with a nonparametric prior (FreeP) (Figure 3A, bottom). Such priors develop a sharp peak around 0.5 ( Figure 12), demonstrating how the prior directly impacts variability.
Finally, for Prediction 4, the close match of the Bayesian best-fitting model to both bias and variability suggests that human behavior is consistent with optimal decoding, whereas the BLO model substantially overestimates this variability and thus the mean square error of the data. Taken together, the Bayesian model offers a principled and accurate account of human behavior in this task.
@@ 4.2 Testing the Generality on Decision-Making Under Risk
All models are formulated within the standard Cumulative Prospect Theory (CPT) framework (Tversky & Kahneman, 1992), in which the subjective utility of a gamble combines a value function $v(x)$ with a probability weighting function $w(p)$. For pure gains,
@@
We first test our framework on the pricing data from Zhang et al. (2020). On each trial, subjects ($N=75$) chose between a two-outcome monetary gamble and a sure amount, with the sequence of choices adaptively converging to the Certainty Equivalent (CE) of the gamble.
All models are expressed within the CPT framework (Eq. 13), differing only in the form of the probability weighting function $w(p)$. BLO treats $w(p)$ as a deterministic transformation of the objective probability, whereas in our Bayesian framework, $w(p)$ arises from stochastic encoding and optimal decoding. Within the Bayesian class, we test uniform, bounded log-odds, and freely fitted encodings, allowing direct comparison both with BLO and across Bayesian variants. Directly fitting CE responses does not isolate the contribution of $w(p)$, since each CE reflects both the utility function and probability weighting. Because BLO defines $\hat{\pi}(p)$ deterministically, it can be directly fit to CE responses, as in Zhang et al. (2020). By contrast, in the Bayesian framework, $\hat{\pi}(p)$ is derived from a stochastic encoding. For clearer inference on $w(p)$, we include an intermediate step: each reported CE is first inverted through the CPT utility function to obtain an implied probability weight for that trial, and then the Bayesian encoding parameters are estimated from these trial-level weights. With these encoding parameters fixed, the model predicts a distribution over CEs for each gamble; we then re-optimize the utility exponent $\alpha$ and CE noise variance $\sigma^{2}_{\text{CE}}$ to maximize the likelihood of the observed CE data. This ensures a fair evaluation that the final model comparison is based on the likelihood of the original CE responses for all models.
@@
Figure 5: Across both pricing (A–B) and choice tasks (C–D), Bayesian models recover U-shaped resources with peaks near 0 and 1, consistent with Prediction 1. In the pricing task, all Bayesian variants outperform BLO, though the fits do not clearly distinguish between U-shaped and uniform encoding (B). In the choice task, however, the freely fitted encoding provides a markedly better fit than a uniform encoding and also outperforms standard parametric weighting functions (C–D).
@@
To further test the generality of our framework, we analyzed a subset of the Choice Prediction Competition 2015 dataset (Erev et al., 2017). Specifically, we focused on the subset of two-outcome gambles with fully described probabilities and uncorrelated payoffs (Amb=0, Corr=0, LotNum=1), yielding 187,150 trials across 153 subjects.
In the choice task, probabilities of gains and losses are modeled with separate weighting functions $w^{+}(p)$ and $w^{-}(p)$, following CPT. The Bayesian model implements these by fitting separate priors for gains and losses while sharing the same encoding, with $w^{\pm}(p)$ computed as the expectation of $\hat{p}$ under the encoding distribution. As parametric benchmarks, we included the LILO function (Eq. 1) and the Prelec function (Prelec, 1998). BLO was not considered here, as its original specification does not extend to choice tasks. Because the observed responses are binary choices rather than CE values, we modeled choice probabilities with a logit rule applied to the difference in subjective utilities, with a temperature parameter $\tau$. All parameters were estimated separately for each subject.
As shown in Figure 5D, the Bayesian model with freely fitted prior and encoding fits the data better than both Prelec and LILO, demonstrating the framework’s generality to the choice task. The fitted resources are U-shaped, with high sensitivity near 0 and 1 (Figure 5C). Crucially, the freely fitted encoding performs significantly better than a model constrained to a uniform encoding. This finding confirms Prediction 1 and provides strong evidence that optimal decoding from U-shaped resources generalizes to decision-making under risk.
@@ 4.3 Testing Bayesian Account via Adaptation to Stimulus Statistics
A central prediction of Bayesian accounts is that priors adapt to the statistics of the environment. In our setting, this implies that a bimodal stimulus distribution should induce a bimodal prior, yielding biases attracted to the closest mode (Prediction 2). BLO, by construction, is limited to a single anchor and therefore can only predict one point of attraction. Efficient coding with matched prior and encoding (Frydman & Jin, 2023) makes yet another distinct prediction: biases away from the prior modes (Wei & Stocker, 2015).
@@
The bimodal stimulus distribution induced a distinctive bias pattern in human responses, with two cross-over points centered on the distribution’s modes, indicating bias towards the modes (Figure 6A). Bayesian models reproduced this pattern: the FreeP+FreeE variant not only captured the observed bias (B) but also recovered a bimodal prior with peaks aligned to the stimulus modes (C). Quantitatively, Bayesian models outperformed BLO (D), with BimodalP providing a better fit than a unimodal Gaussian prior, and FreeP achieving the best overall performance. In contrast, BLO failed to reproduce the multi-peaked bias, and the efficient-coding prediction of biases is away from the modes (Appendix Figure 32). Overall, these results confirm Prediction 2: Bayesian framework with the Prior Attraction could produce deviation from standard S-shaped bias.
@@ 5 Discussion and Conclusion
Our results suggest that human probability distortion can be parsimoniously explained as optimal decoding from noisy neural encodings. Across tasks (dot counting, lottery pricing, lottery choice), we consistently find U-shaped encoding resources with peaks near 0 and 1. This nonuniformity induces systematic biases toward the interior of $[0,1]$, providing a principled account of the classic S-shaped weighting function.Alternative explanations—such as regression away from response boundaries (Fennell & Baddeley, 2012; Bedi et al., 2025) or efficient coding with fully matched priors and encodings (Frydman & Jin, 2023)—fit the data less well. Our findings do not rule out efficient coding but suggest that priors and encodings may adapt on different timescales, preventing a full match of prior and encoding (Fritsche et al., 2020).
@@
In conclusion, this work adds to a growing line of research linking decision-making distortions to imprecise yet structured mental representations (e.g. Woodford, 2012; Khaw et al., 2021; Frydman & Jin, 2022; Barretto-García et al., 2023; Zhang et al., 2020; Frydman & Jin, 2023). By grounding probability representation in Bayesian decoding of noisy encodings, our framework both unifies prior accounts and improves quantitative fit over the state-of-the-art Zhang et al. (2020) across both perceptual judgement and decision-making under risk.
@@
Complete derivations and proofs for the theoretical results are provided in Appendix B. All datasets used in Sections 4.1 and 4.2 are publicly available. Any minimal preprocessing is described in Appendix C.1. For modeling, we build on the publicly released code of (Hahn & Wei, 2024). Our scripts to reproduce all figures and experiments will be available upon publication.
Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. X.-X. Wei is supported by a Sloan Research Fellowship from the Alfred P. Sloan Foundation. The authors thank Yash Sarrof and Entang Wang for feedback on the paper.
@@ B.1.1 Properties of Bayesian Model
At any $p\in(0,1)$, the Bayesian model has the bias, assuming $F(0),F(1)$ are finite:
$\underbrace{\operatorname{sign}(0.5-p)\frac{A_{1,\sigma}(p)}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{(1-A_{2,\sigma}(p))\left(\frac{1}{\mathcal{J}(p)}\right)^{\prime}}_{\text{Likelihood Repulsion}}+\underbrace{(1-A_{3,\sigma}(p))\frac{1}{\mathcal{J}(p)}\left(\log P_{{prior}}(p)\right)^{\prime}}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$ (14)
as $\sigma\rightarrow 0$. where $A_{\dots,\sigma}(p)$ are bounded, positive, and converge to 0 as $\min\{|F(p)-F(0)|/\sigma,|F(p)-F(1)|/\sigma\}$ increases. Second, if $F(0)=$-$\infty$, $F(1)=\infty$, the theorem remains valid with $A_{\dots,\sigma}(p)\equiv 0$.
@@
The case were $F(0)=$-$\infty$, $F(1)=\infty$ is immediate from Theorem 1 in Hahn & Wei (2024). The case where $F(0),F(1)$ are finite is an adaptation of Theorem 3 in Hahn & Wei (2024), which examines the case where the stimulus space has one boundary (e.g., $(-\infty,\theta_{Max}]$). For this case, it proves the decomposition (we consider the special case of $p=2$, as we assume the posterior mean estimator):
@@
where $D:=\frac{F(\theta_{Max}-F(\theta)}{\sigma}$, and where $C_{\dots}$ are all positive, $C_{1,\dots}=\Theta(1)$ as $\sigma\rightarrow 0$, $C_{1,\dots}\rightarrow_{D\rightarrow\infty}0$, $C_{2/3,\dots}\rightarrow_{D\rightarrow\infty}1$. In fact, $C_{3,p,D},C_{2,p,D}\in[0,1]$. Now we partition $(0,1)$ into $(0,1/2)$ and $(1/2,1)$; for the second part, we use the above decomposition, for the first part, we use a mirror-image where the stimulus space is $0,\infty)$. By appropriately defining $A_{1,\sigma},A_{2,\sigma},A_{3,\sigma}$, we obtain
$\underbrace{\operatorname{sign}(0.5-p)\frac{A_{1,\sigma}(p)}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{(1-A_{2,\sigma}(p))\left(\frac{1}{\mathcal{J}(p)}\right)^{\prime}}_{\text{Likelihood Repulsion}}+\underbrace{(1-A_{3,\sigma}(p))\frac{1}{\mathcal{J}(p)}\left(\log P_{{prior}}(p)\right)^{\prime}}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$ (16)
@@
The term $\left[\mathbb{E}[\hat{p}|p]-\mathbb{E}[F^{-1}(m)|p]\right]$ corresponds to the quantity referred to as Decoding Bias in Hahn & Wei (2024); Hahn et al. (2025). The equation is then obtained from Lemma S24 in Hahn et al. (2025). Note that, for the quantity $C_{dec,M}$ used in that lemma, we have $\left[\mathbb{E}[\hat{p}|p]-\mathbb{E}[m|p]\right]=\sigma^{2}C_{dec,M}+\mathcal{O}(\sigma^{4})$, completing the proof. ∎

### v2

@@ 4 Empirical Validation of the Bayesian Framework
We empirically evaluate our theory within a unified Bayesian framework. To test our model and predictions against the full landscape of existing theories, we perform a head-to-head comparison of our Bayesian account (U-shaped FI with task-dependent priors) against: i) Fixed probability distortion function, ii) BLO model (zhang2020), iii) Bayesian model with uniform encoding (fennell2012uncertainty; bedi2025probability), and iv) Efficient coding with matched prior and encoding (frydman2023source). This comparison allows us to pinpoint the specific limitations of each account.
First, we validate Predictions 1, 3, 4 using a judgment of relative frequency (JRF) task, where subjects estimate the proportion of dots. Here, we show that uniform encoding models cannot capture the S-shaped bias patterns, while BLO cannot account for response variability. Second, we assess the generality of our framework by applying it to a different domain, decision-making under risk, showing that our account outperforms classical fixed distortion functions and other accounts. Third, we test our model’s ability to adapt to different stimulus statistics to test Prediction 2 and demonstrate that BLO cannot explain human adaptation to bimodal distributions.
@@ 4.1 Validating Predictions on Judgment of Relative Frequency (JRF) data
We analyze data from zhang2020 and ZM2012, where subjects ($N=86$) judge the percentage of dots of a target color in arrays of black and white dots on a gray background. [^1] A complete description of the tasks and datasets used in this paper is available in Appendix C.2.
Given that BLO outperforms traditional fixed distortion functions (zhang2020), we treat it as the strong baseline for parametric models. We further investigate whether our Bayesian account offers a more unified explanation than competing accounts of uniform encoding and efficient coding. We evaluated Bayesian variants combining Gaussian prior with fitted mean and variance (GaussianP) or nonparametric prior(FreeP) with three encodings: uniform encoding, $F(p)=p$ (UniformE), bounded-log-odds encoding (BoundedLOE) [^2] Formulated as $F(p)=\log(p+\beta_{1})-\log(1-p+\beta_{2})$, similar to11 but with separate upper and lower bounds, in line with BLO’s $\Delta_{-},\Delta_{+}$ bounds., and freely fitted $F$ (FreeE). We also included a matched prior–encoding model to test the efficient coding account. In both BLO and the Bayesian model, parameters are optimized for each subject individually by maximizing trial-by-trial data likelihood. We fitted the Bayesian model using the method of hahn2024unifying, which permits identifying the encoding nonparametrically from behavioral responses (Hahn2025Identifiability). [^3] Across all datasets, we ran all models using a grid of 200 points and regularization strength 1.0, ensuring a head-to-head comparison.
@@
First, we examine the shape of resources (Prediction 1). Models with nonuniform encoding capture the characteristic S-shaped bias in behavior, whereas the variant with uniform encoding fails to do so (Figure 3A, top). To pinpoint the mechanism, we decomposed the bias (Figure 4). The UniformE model exhibits zero Likelihood Repulsion, forcing it to rely exclusively on Boundary Regression and Prior Attraction, which are insufficient to recover the full S-shaped bias. In contrast, the recovered nonuniform resources are consistently U-shaped with peaks near the boundaries (Figure 4), a pattern that generates the required repulsion. While the BLO model also shares the same overall U-shaped resources, it has two discontinuities, a property proved in Appendix Theorem 8 and illustrated in Appendix Figure 11.Furthermore, contrary to the efficient encoding, which predicts that encoding matches the prior, we found that the recovered resources and prior have distinct shapes (Appendix Figure 32). Consequently, enforcing a matched prior-encoding constraint yields a substantially inferior fit compared to decoupled fitting of both components (Appendix Figure 8).
Figure 4: Resources (A) and bias decomposition (B) of FreeP FreeE and FreeP UniformE models for JRF data. IQR denotes interquartile range across subjects. In the best-fitting FreeP FreeE model, boundary regression plays only a small role, whereas Likelihood Repulsion consistently produces a bias away from 0 and 1.
Next, we test the fit of response variability (Prediction 3). The bimodal pattern of response variability with a dip near 0.5 is only explained by models with a nonparametric prior (FreeP) (Figure 3A, bottom). Such priors develop a sharp peak around 0.5 (Figure 12) which mathematically suppresses variance. In contrast, The BLO model predicts variability based on noise in log-odds space, typically implying a simple inverted-U shape (Theorem 10), failing to capture this structure.
Finally, for Prediction 4, the close match of the Bayesian best-fitting model to both bias and variability suggests that human behavior is consistent with optimal decoding, whereas the BLO model substantially overestimates this variability and thus the mean square error of the data. Taken together, these empirical results validate prediction 1, 3 and 4, showing that probability distortion stems from optimal decoding of U-shaped neural representations rather than from uniform encoding or deterministic transformations as in BLO.
@@ 4.2 Testing the Generality on Decision-Making Under Risk
All models are formulated within the standard Cumulative Prospect Theory (CPT) framework (tversky1992advances), in which the subjective utility of a gamble combines a value function $v(x)$ with a probability weighting function $w(p)$. For pure gains,
@@
We first test our framework on the pricing data from zhang2020. On each trial, subjects ($N=75$) chose between a two-outcome monetary gamble and a sure amount, with the sequence of choices adaptively converging to the Certainty Equivalent (CE) of the gamble.
Given that  zhang2020 showed BLO as a superior account compared to traditional functions in this pricing task as well, it serves as the primary benchmark of parametric models. All models are expressed within the CPT framework (Eq. 13), differing only in the form of the probability weighting function $w(p)$. BLO treats $w(p)$ as a deterministic transformation of the objective probability, whereas in our Bayesian framework, $w(p)$ arises from stochastic encoding and optimal decoding. Within the Bayesian class, we test uniform, bounded log-odds, and freely fitted encodings, allowing direct comparison both with BLO and across Bayesian variants. Directly fitting CE responses does not isolate the contribution of $w(p)$, since each CE reflects both the utility function and probability weighting. Because BLO defines $\hat{\pi}(p)$ deterministically, it can be directly fit to CE responses, as in zhang2020. By contrast, in the Bayesian framework, $\hat{\pi}(p)$ is derived from a stochastic encoding. For clearer inference on $w(p)$, we include an intermediate step: each reported CE is first inverted through the CPT utility function to obtain an implied probability weight for that trial, and then the Bayesian encoding parameters are estimated from these trial-level weights. With these encoding parameters fixed, the model predicts a distribution over CEs for each gamble; we then re-optimize the utility exponent $\alpha$ and CE noise variance $\sigma^{2}_{\text{CE}}$ to maximize the likelihood of the observed CE data. This ensures a fair evaluation that the final model comparison is based on the likelihood of the original CE responses for all models.
@@
Figure 5: Across both pricing (A–B) and choice tasks (C–D), Bayesian models recover U-shaped resources with peaks near 0 and 1, consistent with Prediction 1. In the pricing task, all Bayesian variants outperform BLO, though the fit do not clearly distinguish between U-shaped and uniform encoding (B). In the choice task, however, the freely fitted encoding provides a markedly better fit than a uniform encoding and also outperforms standard parametric weighting functions (C–D).
@@
To further test the generality of our framework, we analyzed a subset of the Choice Prediction Competition 2015 dataset (erev2017anomalies). Specifically, we focused on the subset of two-outcome gambles with fully described probabilities and uncorrelated payoffs (Amb=0, Corr=0, LotNum=1), yielding 187,150 trials across 153 subjects.
In the choice task, probabilities of gains and losses are modeled with separate weighting functions $w^{+}(p)$ and $w^{-}(p)$, following CPT. The Bayesian model implements these by fitting separate priors for gains and losses while sharing the same encoding, with $w^{\pm}(p)$ computed as the expectation of $\hat{p}$ under the encoding distribution. As parametric benchmarks, we included the LILO function (Eq. 1) and the Prelec function (prelec1998probability). BLO was not considered here, as its original specification does not extend to choice tasks. Because the observed responses are binary choices rather than CE values, we modeled choice probabilities with a logit rule applied to the difference in subjective utilities, with a temperature parameter $\tau$. All parameters were estimated separately for each subject.
As shown in Figure 5D, the Bayesian model with freely fitted prior and encoding fits the data better than both Prelec and LILO, demonstrating the framework’s generality to the choice task. The fitted resources are U-shaped, with high sensitivity near 0 and 1 (Figure 5C). Crucially, the freely fitted encoding performs significantly better than a model constrained to a uniform encoding. This finding confirms Prediction 1 and provides strong evidence that optimal decoding from U-shaped resources generalizes to decision-making under risk. In sum, these findings extend the validity of Prediction 1 to the economic domain, showing that U-shaped resources are essential for capturing the discriminability required in risky choice, whereas uniform encoding is structurally insufficient.
@@ 4.3 Testing Bayesian Account via Adaptation to Stimulus Statistics
A central prediction of Bayesian accounts is that priors adapt to the statistics of the environment. In our setting, this implies that a bimodal stimulus distribution should induce a bimodal prior, yielding biases with attraction points near each mode (Prediction 2). BLO, by construction, is limited to a single anchor and therefore can only predict one point of attraction. Efficient coding with matched prior and encoding (frydman2023source) predicts that resources reallocate to match the bimodal prior, which would theoretically generate a repulsion bias away from the modes (wei2015bayesian). In sharp contrast, our account predicts attraction towards the modes of the prior.
@@
The bimodal stimulus distribution induced a distinctive bias pattern in human responses, with two cross-over points centered on the distribution’s modes (Figure 6A). Bayesian models reproduced this pattern: the FreeP+FreeE variant not only captured the observed bias (B) but also recovered a bimodal prior with peaks aligned to the stimulus modes (C). Quantitatively, Bayesian models outperformed the matched prior-encoding model(efficient coding account) and BLO (D), with BimodalP providing a better fit than a unimodal Gaussian prior, and FreeP achieving the best overall performance. In contrast, BLO predicts a static S-shape and fails to reproduce the multi-peaked structure. Furthermore, the Efficient Coding model predicts repulsion away from the modes (Appendix Figure 31), a pattern diametrically opposed to the attraction observed in human data. Overall, these results confirm Prediction 2 and demonstrate that human probability perception involves a flexible prior distinct from resource allocation.
@@ 5 Discussion and Conclusion
Our results suggest that human probability distortion can be parsimoniously explained as optimal decoding from noisy neural encodings. Across tasks (dot counting, lottery pricing, lottery choice), we consistently find U-shaped encoding resources with peaks near 0 and 1. As we show in both theory and modeling, this nonuniformity induces systematic biases toward the interior of $[0,1]$, providing a principled account of the classic S-shaped weighting function. Importantly, we find this account not only provides better fit than fixed probability weighting and the Bayesian Log-Odds Models, but also better fits than competing Bayesian accounts. Alternative explanations – such as regression away from response boundaries (fennell2012uncertainty; bedi2025probability) or efficient coding with fully matched priors and encodings (frydman2023source) – fit the data less well. Our findings do not rule out efficient coding but suggest that priors and encodings may adapt on different timescales, preventing a full match of prior and encoding (fritsche2020bayesian).
@@
In conclusion, this work adds to a growing line of research linking decision-making distortions to imprecise yet structured mental representations (e.g. woodford2012prospect; khaw2021cognitive; frydman2022efficient; barretto2023individual; zhang2020; frydman2023source). By grounding probability representation in Bayesian decoding of noisy encodings, our framework both unifies prior accounts and improves quantitative fit over the state-of-the-art zhang2020 across both perceptual judgement and decision-making under risk.
@@
Complete derivations and proofs for the theoretical results are provided in Appendix B. All datasets used in Sections 4.1 and 4.2 are publicly available. Any preprocessing is described in Appendix C.2.
For modeling, we developed an implementation extending the optimization framework of hahn2024unifying. Specific adaptations required for probability domain and task-specific processes are detailed in Appendix C.6C.8C.9.2. The reimplementation of zhang2020 is detailed in Appendix C.11C.12. Our scripts to reproduce all figures and experiments will be available upon publication.
@@ B.1.1 Properties of Bayesian Model
There are functions $A_{1,\sigma},A_{2,\sigma},A_{3,\sigma}:[0,1]\rightarrow\mathbb{R}$ that are bounded, positive, and satisfy
$A_{\dots,\sigma}(p)\leq\phi\left(\frac{\sigma}{\min\{|F(p)-F(0)|,|F(p)-F(1)|\}}\right)$ (14)
for some nondecreasing function $\phi:\mathbb{R}\rightarrow\mathbb{R}$ with $\lim_{x\downarrow 0}\phi(x)=0$, across $p\in(0,1);\sigma>0$. Then, at any $p\in(0,1)$, the Bayesian model has the bias, as $\sigma\rightarrow 0$:
$\operatorname{Bias}(p)=\underbrace{A_{1,\sigma}(p)\cdot\frac{\operatorname{sign}(0.5-p)}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{(1-A_{2,\sigma}(p))\cdot\frac{\operatorname{d}}{\operatorname{d}p}\left(\frac{1}{\mathcal{J}(p)}\right)}_{\text{Likelihood Repulsion}}$
$+\underbrace{(1-A_{3,\sigma}(p))\frac{1}{\mathcal{J}(p)}\cdot\frac{\operatorname{d}}{\operatorname{d}p}\left(\log P_{{prior}}(p)\right)}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$
@@
We first consider the case where $F(0)=$-$\infty$, $F(1)=\infty$, in which case (14) forces $A_{\dots,\sigma}(p)\equiv 0$. The above expression with $A_{\dots,\sigma}(p)\equiv 0$ simplifies to:
$\underbrace{\frac{\operatorname{d}}{\operatorname{d}p}\left(\frac{1}{\mathcal{J}(p)}\right)}_{\text{Likelihood Repulsion}}+\underbrace{\frac{1}{\mathcal{J}(p)}\cdot\frac{\operatorname{d}}{\operatorname{d}p}\left(\log P_{{prior}}(p)\right)}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$
Indeed, this is exactly the expression from Theorem 1 in hahn2024unifying; the claim thus follows from that theorem.
We next consider the case where $F(0),F(1)$ are finite; we obtain this on the basis of Theorem 3 in hahn2024unifying. That theorem examines the bias in the setting where the stimulus space has one boundary (e.g., $(-\infty,\theta_{Max}]$). For this case, it proves the decomposition (we consider the special case of $p=2$, as we assume the posterior mean estimator):
@@
where $D(p):=\frac{F(\theta_{Max})-F(p)}{\sigma}$, and where $C_{\dots}$ are given as follows:
  1. 1.
    $C_{1,p,D,F,\sigma}$ depends on the probability $p$, the quantity $D(p)$, the encoding function $F$, and the sensory noise magnitude $\sigma$
  2. 2.
    $C_{2,p,D}$ depends on the probability $p$ and the quantity $D(p)$
  3. 3.
    $C_{3,p,D}$ depends on the probability $p$ and the quantity $D(p)$
  4. 4.
    $C_{1,p,D,F,\sigma}$, $C_{2,p,D}$, $C_{3,p,D}$ are positive everywhere
  5. 5.
    $C_{1,p,D,F,\sigma}=\Theta(1)$ when $\sigma\rightarrow 0$
  6. 6.
    $C_{1,\dots}\rightarrow_{D\rightarrow\infty}0$, $C_{2/3,\dots}\rightarrow_{D\rightarrow\infty}1$.
In fact, inspecting the proof of the theorem shows that $C_{3,p,D},C_{2,p,D}\in(0,1)$ for all values of $p,D$, and that $C_{1,p,D(p),F,\sigma}$ is continuous in $p$ and $\sigma$ for $\sigma>0$.
Our aim is to use this result to establish the bias in the setting where the stimulus space is $[0,1]$, i.e., has two boundaries.
Towards this end, we partition $(0,1)$ into $(0,1/2)$ and $(1/2,1)$.
First, $p\in(1/2,1)$, the above decomposition is valid with $\theta_{Max}=1$. The reason is because the contribution of the lower boundary at 0 to the bias can be absorbed into the $\mathcal{O}(\sigma^{4})$ remainder, because only exponentially small probability mass of the likelihood of the encoding $m$ falls into regions outside a small local environment of $p$ (hahn2024unifying, SI Appendix, Section S.3.1.1).
Second, for $p\in(0,1/2)$, by transforming the above result via $p\mapsto-p$, we obtain for the stimulus space $\theta_{Min},\infty)$ the bias:
$\underbrace{\frac{C_{1,p,D,F,\sigma}}{\sqrt{\mathcal{J}(p)}}}_{\text{Regression from Boundary}}+\underbrace{C_{3,p,D}\left(\frac{1}{\mathcal{J}(p)}\right)^{\prime}}_{\text{Likelihood Repulsion}}+\underbrace{C_{2,p,D}\frac{1}{\mathcal{J}(p)}\left(\log P_{{prior}}(p)\right)^{\prime}}_{\text{Prior Attraction}}+\mathcal{O}(\sigma^{4})$ (16)
where $D(p):=\frac{F(p)-F(\theta_{Min})}{\sigma}$, and where $C_{\dots}$ satisfy exactly the same conditions as given above.
Now write $\hat{C}_{...}$ for the coefficients for $p\in(1/2,1)$ and $\tilde{C}_{...}$ for the coefficients for $p\in(0,1/2)$; the same with $\hat{D}(p)$ and $\tilde{D}(p)$.
Now define
$A_{1,\sigma}(p)$ $:=\hat{C}_{1,p,D(p),F,\sigma}\cdot 1_{p\geq 1/2}+\tilde{C}_{1,p,D(p),F,\sigma}\cdot 1_{p<1/2}$
$A_{2,\sigma}(p):=$ $1-\left(\hat{C}_{3,p,D(p)}\cdot 1_{p\geq 1/2}+\tilde{C}_{3,p,D(p)}\cdot 1_{p<1/2}\right)$
$A_{3,\sigma}(p):=$ $1-\left(\hat{C}_{2,p,D(p)}\cdot 1_{p\geq 1/2}+\tilde{C}_{2,p,D(p)}\cdot 1_{p<1/2}\right)$
First, inserting these definitions shows that the two decompositions do indeed produce the one claimed by the theorem.
We now need to show the remaining conclusions. By the assumptions above, $A_{1,\sigma}$ is positive and also continuous in both $\sigma$ and $p$. Next, by the facts that $C_{1,p,D,F,\sigma}=\Theta(1)$ when $\sigma\rightarrow 0$ and $C_{1,\dots}\rightarrow_{D\rightarrow\infty}0$, we find that $A_{1,\sigma}(p)$ is bounded across $\sigma$ and $p$ whenever $\sigma$ is upper-bounded by some $\sigma_{Max}$. As $C_{2,\dots},C_{3,\dots}\in[0,1]$, we find that $A_{2,\sigma}$ and $A_{3,\sigma}$ are positive and bounded. We have thus shown that $A_{1,\sigma}$, $A_{2,\sigma}$, $A_{3,\sigma}$ are positive and bounded.
Finally, as mentioned above, $\hat{C}_{1,p,\hat{D}(p),F,\sigma}\rightarrow 0$ as $\hat{D}(p)\rightarrow\infty$; equivalently,
$\hat{C}_{1,p,\hat{D}(p),F,\sigma}=\mathcal{O}\left(\phi_{1}\left(\frac{1}{\hat{D}(p)}\right)\right)$ (17)
for some nondecreasing function $\phi_{1}$ with $\lim_{s\downarrow 0}\phi_{1}(s)=0$. The same holds for $\tilde{C}$, $\tilde{D}$ with a function $\phi_{2}$. Thus,
$A_{1,\sigma}(p)=$ $\mathcal{O}\left(\phi_{1}\left(\frac{1}{\hat{D}(p)}\right)\right)+\mathcal{O}\left(\phi_{2}\left(\frac{1}{\tilde{D}(p)}\right)\right)$
$=$ $\mathcal{O}\left(\max\left\{\phi_{1}(\frac{1}{\hat{D}(p)}),\phi_{2}(\frac{1}{\tilde{D}(p)})\right\}\right)$
which is equivalent to the claim that
$A_{1,\sigma}(p)\leq\phi\left(\frac{\sigma}{\min\{|F(p)-F(0)|,|F(p)-F(1)|\}}\right)$ (18)
across $p\in(0,1);\sigma>0$, for some nondecreasing function $\phi$ with $\lim_{s\downarrow 0}\phi(s)=0$. An analogous argument applies to $A_{2,\sigma}$ and $A_{3,\sigma}$.
@@
The term $\left[\mathbb{E}[\hat{p}|p]-\mathbb{E}[F^{-1}(m)|p]\right]$ corresponds to the quantity referred to as Decoding Bias in hahn2024unifying; Hahn2025Identifiability. The equation is then obtained from Lemma S24 in Hahn2025Identifiability. Note that, for the quantity $C_{dec,M}$ used in that lemma, we have $\left[\mathbb{E}[\hat{p}|p]-\mathbb{E}[m|p]\right]=\sigma^{2}C_{dec,M}+\mathcal{O}(\sigma^{4})$, completing the proof. ∎